# Automatically Detecting Checked-In Secrets in Android Apps: How Far Are We?

### v1

@@ 3.2. RQ1-Result
For the representative static analysis tool, we chose LeakScope because it is the only static analysis tool designed for Android APK files and is available on GitHub (Lea, 2020).
For machine learning-based tools, we selected PassFinder for further evaluation. PassFinder, SecretHunter and Seagull are the only publicly available tools. Among them, SecretHunter aims to address the issue of missing files in GitHub repositories during the scanning process, often due to API and bandwidth limitations. Since SecretHunter is tailored specifically for the GitHub platform, we excluded it from our study. As for Seagull, it highly relies on the results provided by CodeQL (cod, 2024), , which is incompatible with our focus on Android apps. Furthermore, the key idea of Seagull is very similar to PassFinder, which takes contextual information (i.e., the surrounding code snippets) as input for the models. Given their similarity in leveraging context for detection, we selected PassFinder as the representative tool for machine-learning-based analysis in this study.
@@
Three-Layer Filter. Meli et al. (Meli et al., 2019) is the first to conduct a large-scale analysis on the checked-in secret problem of the open-source projects on GitHub, covering 13% of all the public repositories. Their methodology involved initially extracting strings from the repositories through regular expression matching. Subsequently, these candidate strings go through a Three-Layer Filter by evaluating the candidate string’s entropy value, pattern, and word. First, the authors designed 15 regular expressions for 12 different cloud providers and applied the 15 regular expressions to extract candidate strings from the source code. Then, the extracted candidate strings will be filtered by three layers. We refer to this technique as Three-Layer Filter for the rest of this paper. Specifically, the three filers are:
- Entropy Filter. The entropy filter would first calculate the Shannon Entropy (Shannon, 1951) for each string, and within the same regular expression group, a string will fail if its Shannon entropy is three standard deviations away from the mean.
@@
LeakScope (Zuo et al., 2019) is the first automated tool that aims to detect the data leakage problem caused by checked-in secret leakage and misconfiguration of secret permission in Android apps. Their tool is developed based on the key finding that if a project wants to use the services provided by cloud providers, the project would likely call the cloud APIs provided by the cloud providers (i.e., the APIs defined in the cloud SDKs). These cloud APIs generally take one or several checked-in secrets as part of their parameters. Take the CloudStorageAccount.parse(String) API of Microsoft Azure as an example; this cloud API takes a checked-in secret as a parameter and returns the account that is connected to this checked-in secret (clo, 2024); LeakScope collected 32 cloud APIs that take checked-in secret as parameter from three cloud providers: Microsoft Azure, Amazon AWS and Google Firebase. To extract the value of the checked-in secret, LeakScope performs an inter-procedural backward analysis starting from the cloud APIs; along with the backward analysis, LeakScope keeps track of the data flow associated with the parameter of this cloud API. Through this data flow tracking, LeakScope proceeds to compute the value of the parameter. Moreover, LeakScope employs a signature generation mechanism for cloud APIs based on their package structure and content, enabling identification of their usage within the project. This signature empowers LeakScope to operate effectively, even in scenarios where APK files are obfuscated.
PassFinder (Feng et al., 2022) is a checked-in secret detection tool utilizing the deep neural network. It conducts an analysis of strings by considering both the intrinsic value of the string itself and its contextual surroundings, particularly encompassing twelve lines of context. PassFinder trains separate models for the intrinsic value and context of the string. The prediction is given by combining the results of two models. PassFinder is designed to analyze the password and API credentials in the GitHub repository as it is trained with the code snippet collected from the GitHub repository. It is also capable of handling different programming languages.
4. RQ2: Comparsion of Exisiting TOOLs
@@
4.1. RQ2-Methodology
We aim to compare the performance of the selected tools by answering the following two sub-research questions:
  • RQ2.1 (Number of detected checked-in secrets): How many checked-in secrets can be detected by each of the selected tools?
  • RQ2.2 (False positives): How many checked-in secrets reported by each tool are false positives?
  • RQ2.3 (False negatives): How many false negatives are produced by each tool?
4.1.1. Dataset
We first collected a dataset of Android apps as the evaluation subjects. We leveraged Raccoon APK Downloader (rac, 2024) to download Android apps from Google Play Store. We selected the top 200 Android apps in each category based on the number of downloads according to the statistics on AppBrain (app, 2024). After excluding the Android apps that failed to be processed by Soot, we collected 5,135 APK files in total.
@@ 4.1.1. Dataset
4.1.2. Tool Adaptation
None of our selected tools are ready to use for our experiments. The Three-Layer Filter did not provide the necessary code, LeakScope did not provide the cloud API signatures, and PassFinder did not provide the pre-trained models. Thus, some implementation efforts were necessary for our experiments. Additionally, unlike PassFinder, whose detection capability is not strictly constrained, LeakScope and the Three-Layer Filter can only detected the checked-in secrets that are used as inputs in the pre-defined cloud APIs or adhere to the pre-defined regular expressions. Therefore, expanding the list of cloud APIs and regular expressions was necessary to equalize the detection capabilities of LeakScope and the Three-Layer Filter and conduct a fair comparison.
@@ 4.1.2. Tool Adaptation
Adaptation for Three-Layer Filter. As the source code for the filtering process of the Three-Layer Filter is not provided by the authors. We first implemented the necessary code for the filtering process following the explanation of author (Meli et al., 2019). For the string extraction, the source code is not directly accessible in APK files, we used Soot (soo, 2024), which is capable of identifying all the variables of type java.lang.String to extract all the strings within the APK files.
The authors provide regular expressions. However, we noticed that in their experiment, their filtering process only excluded 2% of the strings due to their precise regular expressions. Their regular expression rarely mismatch checked-in secrets. As a result, the filtering process has minimal impact on reducing false positives in their experiment.
However, it is challenging to obtain precise regular expressions for all the categories of checked-in secrets in practice. It is more common for cloud providers to adopt loose regular expressions. For example, checked-in secrets of Twitter and Facebook do not exhibit a strong pattern. Any combination of letters and numbers that fall within a specific range of lengths can match the regular expressions. Loose regular expressions would match more non-secret strings compared to precise regular expressions. Passing more non-secret strings to the filtering process is necessary to fully validate the performance of the Three-Layer Filter.
Therefore, we derived two more regular expressions, as shown in Table 3. We chose Twitter client ID and secret because the corresponding checked-in secrets do not exhibit a strong pattern; Twitter APIs are commonly used in Android apps.
Furthermore, the authors do not provide the source code for the Three-Layer Filter. Therefore, we also implemented the necessary code based on the logic proposed by the authors. The implementation and the full list of collected regular expressions can be found on our website (dat, [n. d.]).
Table 3. Newly proposed regular expressions
@@
Adaptation for LeakScope. In order to use LeakScope to extract the checked-in secrets from an APK file, we have to provide the complete signature of the cloud API and specify which API parameter holds checked-in secrets to LeakScope. The authors provided the names of 32 cloud APIs from Amazon AWS, Microsoft Azure, and Google Firebase, but they did not provide the complete signatures of cloud APIs. Therefore, we also collected the detailed signatures of these 32 cloud APIs. The full list of signatures can be found on our website (dat, [n. d.]).
@@
However, the provided cloud APIs are limited to three cloud providers. The number of supported cloud providers is very limited compared to the Three-Layer Filter. To perform a more in-depth and fair analysis, we aim to collect more cloud APIs that can be analysis starting points for LeakScope, i.e., the cloud APIs that take checked-in secret as a parameter. Given the abundance of cloud providers in the contemporary market, attempting to collect all the cloud APIs that take checked-in secret as parameter from every single provider would be infeasible, especially since some cloud providers may not offer comprehensive API documentation. Therefore, instead of solely relying on the API documentation provided by every cloud provider, our cloud API collection strategy also takes into account the potential cloud API extracted from our dataset.
@@
We first collected 15 cloud providers, including those covered by the Three-Layer Filter but not by LeakScope, as well as the top cloud service providers in 2024 (Slingerland, [n. d.]). Then, we read the available API documentation of the collected cloud providers to identify the cloud APIs that use secrets as part of their parameters. To mitigate the risk of missing some relevant API documentation, we also adopted an additional step. We collected information of all the methods that are called at least once in the Android apps of our dataset. We then selected all the methods that contain the names of our collected cloud providers. By analyzing the context and implementation (API documentation if available) of these methods, we determined whether they are our targeted cloud APIs. For example, a constructor, OSSStsTokenCredentialProvider(String, String, String) was not originally found by reading the API documentation of Alibaba, but as its package name contains Alibaba, we were able to identify this cloud API.
@@
Overall, we collected 13 new cloud APIs from five additional providers. For the remaining 10 cloud providers, we failed to recover any cloud APIs that take checked-in secrets as input. This results in a total of 48 cloud APIs from 10 cloud providers. Table 4 summarizes the distribution of these APIs across the different cloud providers. The authors of LeakScope provided 32 cloud APIs from Amazon, Google and Microsoft. By analyzing API documentation, we identified 10 additional cloud APIs from Alibaba, Facebook, Redhat, Twitter and Stripe. Furthermore, we collected 6 new cloud APIs from Alibaba, Facebook, Paypal and IBM Cloud.
@@
Table 4. Collected cloud APIs and cloud providers’ information that are used as the parameters for LeakScope, the signature of each cloud API can be found on our website (dat, [n. d.]).
Cloud providers Number of Original provided Cloud APIs by LeakScope Number of Newly Collected Cloud APIs by API Documentation Number of Newly Collected Cloud APIs by API Extraction Total Number of Collected Cloud APIs
Amazon AWS 26 0 0 26
Alibaba Cloud 0 2 2 4
Microsoft Azure 4 0 0 4
Facebook/Meta 0 2 1 3
Twitter/X 0 2 0 2
Redhat 0 2 0 2
IBM Cloud 0 0 2 2
Google 2 0 0 2
Paypal Braintree 0 0 1 1
Stripe 0 2 0 2
Adaptation for PassFinder. We contacted the author to request the pre-trained models, but they indicated that the project is not currently being maintained, but the necessary dataset and instruction are provided (pas, 2023). As a result, we followed the exact instructions provided by the author to train the two models provided by PassFinder.
For the password model, we leveraged the original training set provided by the authors (pas, 2023). For the context model, the original training set contains 12,000 code snippets for the ten most popular programming languages on GitHub. For each language, there are 600 snippets with checked-in secrets and 600 snippets without checked-in secrets.
Once the password and context models were trained, we utilized the password model to identify potential checked-in secrets within our dataset. Subsequently, we employed Soot (soo, 2024) to extract the contextual surroundings of the identified checked-in secrets, which were then analyzed by the context model for detection. However, instead of using source code, Soot will output the contents of APK files in Jimple format, an intermediate representation of the Java language. Jimple representations are very similar to Java syntax, so we decided to remain with the Jimple format for the extracted contextual surroundings.
Table 5. The type of checked-in secrets that each tool can detect
@@
The Three-Layer Filter Amazon, Microsoft, Facebook, Twitter client ID, Twitter client secret, Twitter access token, Google API key, Stripe, Square, Picatic, PayPal Braintree, Twilio, MailGun, MailChimp
LeakScope Amazon, Microsoft, Facebook, Twitter client ID, Twitter client secret, Google API key, Stripe, Square Redhat, IBM cloud
@@
4.1.3. Comparison of the Detection Capability of Each Tool
@@
After making the necessary adaptation for each tool, the detection capability of each tool is listed in Table 5. Overall, we extended LeakScope’s detection capability to align with the Three-Layer Filter. However, it is infeasible to identify a comprehensive set of cloud API signatures for LeakScope that matched all the regular expressions used by the Three-Layer Filter and vice versa. This difficulty arises because some cloud providers do not issue secrets that follow specific patterns that can be precisely captured by regular expressions (e.g., Alibaba, Redhat and IBM). Furthermore, we failed to find the cloud APIs of the cloud providers that the Three-Layer Filter covered (e.g., MailGun, MailChimp, and Picatic). Although the detection capability of LeakScope and the Three-Layer Filter is not entirely equalized, our extension ensured a more meaningful comparison among the tools. In the following questions, we also show the detailed detection results of each cloud provider for a fair comparison.
4.1.4. Checked-in Secret Validation and Ethical Considerations.
The common challenge for evaluating the checked-in secret detection tool is the lack of the ground truth. To the best of our knowledge, there is not an available benchmark dataset for checked-in secrets in Android apps. A complete validation of a checked-in secret requires us to access the services associated with the checked-in secret, which would raise ethical concerns. As a result, we decide whether a checked-in secret is valid mainly by manually analyzing its format and contextual surroundings, which is a commonly adopted technique for checked-in secret related research  (Basak et al., 2023a, b). For example, in Listing 2, the presence of some functions from Google Cloud Library is an indicator of checked-in secret. Furthermore, two authors were responsible for manually verifying the potential checked-in secrets. In cases of disagreement, discussions were held where each author presented their reasoning and attempted to reach a consensus through mutual persuasion.
{listing}
@@
Code snippet with one checked-in secret and its context. It is extracted from an Android app in our experiments.
{minted}
[linenos, frame=single, obeytabs=true, numbersep=-10pt]java GoogleApiClient GoogleClient = x.GoogleClient; d.b0.c.l.c(GoogleClient); mGoogleApiClient.disconnect(); GoogleSignInOptions.Builder builder = new Builder; builder = builder.requestIdToken(””); builder = builder.requestEmail();
In addition to manual analysis, we tested the validity of checked-in secrets categorized as Google API keys by sending requests to the Google Map service as they are not related to any form of sensitive information. We also ensured that the request sent by us would not cause any significant financial loss; each checked-in secret was only used once. Furthermore, for any detected checked-in secret, we notified the developers of the corresponding Android apps.
4.2. RQ2.1 Number of Detected Checked-in Secrets
4.2.1. RQ2.1-Methodology
@@
In this research question, we aim to evaluate the effectiveness of the tools by evaluating the number of checked-in secrets that can be detected by each tool. As the detection capability of each tool is not identical, we evaluated the number of detected checked-in secrets identified by each tool across various cloud providers to provide a fairer comparison among the tools. The detailed result of the number of detected checked-in secrets is presented in Table 6. Note that the detected checked-in secrets in this table are not necessarily true positives. In this research question, we demonstrate the overall results reported by each tool. We conduct an analysis on the false positives of the tools in Section 4.3.
@@
Table 6. Number of detected checked-in secrets corresponding to the cloud providers of the three selected tools, NA signifies that the checked-in secrets of the corresponding cloud providers are not in the detection capability of the tool.
@@
The Three-Layer Filter detected the highest number of unique checked-in secrets, while PassFinder failed to detect any. PassFinder’s limited effectiveness can be partially attributed to the context model’s underperformance, which was likely impacted by obfuscation. We further conducted a study to investigate the reasons behind such underperformance of PassFinder. Details of this study is discussed in Section 4.5.
We further categorized all the detected checked-in secrets of each tool to the corresponding cloud providers, as shown in Table 6. Among all the cloud service providers that are supported by each tool, only a few are detected in our dataset. Most of the detected checked-in secrets are from Google cloud service and Twitter cloud service, which is expected given Google’s status as one of the most popular cloud providers. Additionally, LeakScope was able to detect a greater number of Google checked-in secrets compared to the Three-Layer Filter.
For Twitter Client ID/Secret. The Three-Layer Filter identified 88,172, whereas LeakScope only identified 10. This is unsurprising as we adopted loose regular expressions for Twitter when applying the Three-Layer Filter. However, on analyzing a sample of the 88,172 checked-in secrets, despite that they are indeed random strings that appear to be checked-in secrets, none of them are truly Twitter checked-in secrets. This issue will be further discussed in RQ2.2.
@@
4.3. RQ2.2-False Positive
4.3.1. RQ2.2-Methodology
@@
In this research question, we aim to compare the false positives reported by each tool. As described in Section 4.1.4, it is impractical and impossible to fully validate all the obtained checked-in secrets. Consequently, we conducted a manual analysis to identify the false positives. Furthermore, since only checked-in secrets belongs to Google API key and Twitter client ID/Secret are detected by at least two tools. We only selected these checked-in secrets for the comparison of false positives. Additionally, PassFinder is excluded from this comparison as it did not detect any checked-in secrets. A further experiment to comprehensively evaluate PassFinder is detailed in Section 4.5.
@@
However, since the Three-Layer Filter detected 88,172 Twitter client IDs/Secrets, manually validating all of these checked-in secrets was unfeasible. To address this, we randomly sampled 383 checked-in secrets from the 88,172, ensuring a 95% confidence level with a 5% margin of error (ran, 2024). Then, we manually analyzed the randomly sampled Twitter client IDs/Secrets and all the Google API keys to identify any false positive as described in Section 4.1.4.
4.3.2. RQ2.2-Result

### v2

@@ 3.3 Tool Selection for Further Evaluation
For the representative Static Analysis tool, we chose LeakScope because it is the only tool designed for Android APK files and is available on GitHub LeakScopeGithub.
For Machine-Learning-Based tools, we selected PassFinder for further evaluation. PassFinder, SecretHunter, and Seagull are the only publicly available tools. Among them, PassFinder aligns best with the scope of our study. SecretHunter is specifically designed to address the challenge of missing files in GitHub repositories, which often results from API or bandwidth limitations, making it tightly coupled with the GitHub platform and less applicable to our Android-based context. Seagull, on the other hand, relies heavily on CodeQL codeql for code representation, which is incompatible with our analysis setup as it is designed only for open-source projects on Github. Additionally, both PassFinder and Seagull incorporate two key types of features in their detection process: intrinsic string properties and contextual surroundings. Since these two categories represent the most widely adopted feature sets in machine learning-based secret detection, PassFinder’s design makes it a representative choice for our evaluation. Considering practical factors such as tool availability, compatibility, and feasibility of integration into our experimental setup, we selected PassFinder as the representative machine-learning tool.
@@
Three-Layer Filter. Meli et al. meli is the first to conduct a large-scale analysis on the checked-in secret problem of the open-source projects on GitHub, covering 13% of all the public repositories. Their methodology involved initially extracting strings from the repositories through regular expression matching. Subsequently, these candidate strings go through a Three-Layer Filter by evaluating the candidate string’s entropy value, pattern, and word. First, the authors designed 15 regular expressions for 12 different cloud providers and applied the 15 regular expressions to extract candidate strings from the source code. Then, the extracted candidate strings will be filtered by three layers. We refer to this technique as Three-Layer Filter for the rest of this paper. Specifically, the three filers are:
- Entropy Filter. The entropy filter first calculates the Shannon Entropy shannon for each string, and within the same regular expression group, a string will fail if its Shannon entropy is three standard deviations away from the mean.
@@
LeakScope LeakScope is the first automated tool that aims to detect the data leakage problem caused by checked-in secret leakage and misconfiguration of secret permission in Android apps. Their tool is developed based on the key finding that if a project wants to use the services provided by cloud providers, the project would likely call the cloud APIs provided by the cloud providers (i.e., the APIs defined in the cloud SDKs). These cloud APIs generally take one or several checked-in secrets as part of their parameters. Take the CloudStorageAccount.parse(String) API of Microsoft Azure as an example; this cloud API takes a checked-in secret as a parameter and returns the account that is connected to this checked-in secret cloudaccount; LeakScope collected 32 cloud APIs that take checked-in secret as parameter from three cloud providers: Microsoft Azure, Amazon AWS, and Google Firebase. To extract the value of the checked-in secret, LeakScope performs an inter-procedural backward analysis starting from the cloud APIs; along with the backward analysis, LeakScope keeps track of the data flow associated with the parameter of this cloud API. Through this data flow tracking, LeakScope proceeds to compute the value of the parameter. Moreover, LeakScope employs a signature generation mechanism for cloud APIs based on their package structure and content, enabling identification of their usage within the project. This signature empowers LeakScope to operate effectively, even in scenarios where APK files are obfuscated.
PassFinder passfinder is a checked-in secret detection tool utilizing the deep neural network. It conducts an analysis of strings by considering both the intrinsic value of the string itself and its contextual surroundings, particularly encompassing twelve lines of context. PassFinder trains separate models for the intrinsic value and context of the string. The prediction is given by combining the results of these two models. PassFinder is designed to analyze passwords and API credentials in GitHub repositories, as it is trained with code snippets collected from GitHub repositories. It is also capable of handling different programming languages.
Both the Three-Layer Filter and PassFinder were originally designed to operate at the source-code level, where raw code and file structures are directly accessible. In contrast, LeakScope relies on Soot to analyze Java programs by converting them into an intermediate representation known as Jimple. Since our study is based on Android APK files, where the original source code is typically unavailable, we adapted our approach accordingly. Specifically, to enable a fair comparison and apply LeakScope in our setting, we reimplemented the Three-Layer Filter and PassFinder using Soot to process APK files and extract the necessary Jimple representation for analysis.
4 RQ2: Comparison of Existing Tools
@@
4.1 RQ2-Methodology
We aim to evaluate and compare the performance of the selected tools based on the following criteria:
  • RQ 2.1: Number of detected checked-in secrets: How many checked-in secrets can be detected by each of the selected tools?
  • RQ 2.2: False positives: How many checked-in secrets reported by each tool are false positives?
  • RQ 2.3: False negatives: How many false negatives are produced by each tool?
4.1.1 Dataset
We collected a dataset of Android apps for evaluation, using Raccoon APK Downloader raccon to retrieve apps from the Google Play Store. Based on AppBrain appBrain statistics, we selected the top 200 apps in the top 30 categories like Business, Movies, and Music. After filtering out apps that could not be processed by Soot, we obtained a total of 5,135 APK files.
@@ 4.1.1 Dataset
4.1.2 Tool Adaptation
None of the selected tools is ready to use for the experiments. The Three-Layer Filter did not provide the necessary code, LeakScope did not provide the cloud API signatures, and PassFinder did not provide the pre-trained models. Thus, some implementation efforts were necessary for the experiments. Additionally, unlike PassFinder, whose detection capability is not strictly constrained, LeakScope and the Three-Layer Filter can only detect the checked-in secrets that are used as inputs in the pre-defined cloud APIs or adhere to the pre-defined regular expressions. Therefore, expanding the list of cloud APIs and regular expressions was necessary to equalize the detection capabilities of LeakScope and the Three-Layer Filter and conduct a fair comparison.
@@ 4.1.2 Tool Adaptation
Adaptation for Three-Layer Filter. Although the source code for the filtering process of the Three-Layer Filter is not provided by the authors, we implemented the necessary components based on the explanations in the paper meli, which were clear and sufficient to guide our reimplementation. For the string extraction, since the source code is not directly accessible in APK files, we used Soot soot to identifying all the string constants within the APK files, aligning with the original focus of the technique.
The authors provide a set of regular expressions to identify potential API keys. In our reimplementation, we used the same set of regular expressions. However, we observed that in their experiment, the filtering process excluded only 2% of the strings, indicating that their regular expressions were highly precise and rarely mismatched actual API keys. As a result, the filtering process had minimal impact on reducing false positives.
However, it is challenging to obtain precise regular expressions for all the categories of checked-in secrets in practice. It is more common for cloud providers to adopt loose regular expressions. For example, checked-in secrets of Twitter and Facebook do not exhibit a strong pattern. Any combination of letters and numbers that fall within a specific range of lengths can match the regular expressions. To fully evaluate the effectiveness of the filtering mechanism, we introduced two additional, more loose regular expressions from a previous study twitterregex to simulate a scenario where the initial candidate set may contain a higher number of false positives. This allowed us to better assess the contribution of the subsequent filtering process.
Therefore, we used two more regular expressions derived by a previous study twitterregex, as shown in Table 3. We chose Twitter client ID and secret because the corresponding checked-in secrets do not exhibit a strong pattern; Twitter APIs are commonly used in Android apps.
Table 3: Newly proposed regular expressions
@@
Adaptation for LeakScope. In order to use LeakScope to extract the checked-in secrets from an APK file, we have to provide the complete signature of the cloud API and specify which API parameter holds checked-in secrets to LeakScope. The authors of LeakScope provided the names of 32 cloud APIs from Amazon AWS, Microsoft Azure, and Google Firebase, but they did not provide the complete signatures of the cloud APIs. To address this, we consulted the official documentation of each corresponding cloud API to collect the full and accurate signatures of these APIs. The full list of signatures can be found on our website data.
@@
However, the provided cloud APIs are limited to three cloud providers. The number of supported cloud providers is very limited compared to the Three-Layer Filter. To perform a more in-depth and fair analysis, we aim to collect more cloud APIs that can be analysis starting points for LeakScope, i.e., the cloud APIs that take checked-in secret as a parameter. Given the abundance of cloud providers in the contemporary market, attempting to collect all the cloud APIs that take checked-in secret as a parameter from every single provider would be infeasible, especially since some cloud providers may not offer comprehensive API documentation. Therefore, instead of solely relying on the API documentation provided by every cloud provider, our cloud API collection strategy also takes into account the potential cloud API extracted from our dataset.
@@
We first collected 15 cloud providers, including those covered by the Three-Layer Filter but not by LeakScope, as well as the top 13 cloud service providers in 2024 topCP. Then, we read the available API documentation of the collected cloud providers to identify the cloud APIs that use secrets as part of their parameters. To mitigate the risk of missing some relevant API documentation, we also adopted an additional step. We collected information of all the methods that are called at least once in the Android apps of the dataset. We then selected all the methods that contain the names of the collected cloud providers. By analyzing the context and implementation (API documentation if available) of these methods, we determined whether they are our targeted cloud APIs. For example, a constructor, OSSStsTokenCredentialProvider(String, String, String) was not originally found by reading the API documentation of Alibaba, but as its package name contains Alibaba, we were able to identify this cloud API.
@@
Overall, we collected 13 new cloud APIs from five additional providers. For the remaining 10 cloud providers, we failed to recover any cloud APIs that take checked-in secrets as input. This results in a total of 48 cloud APIs from 10 cloud providers. The authors of LeakScope provided 32 cloud APIs from Amazon, Google, and Microsoft. By analyzing API documentation, we identified 10 additional cloud APIs from Alibaba, Facebook, Redhat, Twitter, and Stripe. Furthermore, we collected 6 new cloud APIs from Alibaba, Facebook, Paypal, and IBM Cloud.
@@
Adaptation for PassFinder. We contacted the author to request the pre-trained models, but they indicated that the project is not currently being maintained, but the necessary dataset and instructions are provided passFdATA. As a result, we followed the exact instructions provided by the author to train the two models provided by PassFinder.
For the password model, we leveraged the original training set provided by the authors passFdATA. For the context model, the original training set contains 12,000 code snippets for the ten most popular programming languages on GitHub. For each language, there are 600 snippets with checked-in secrets and 600 snippets without checked-in secrets.
Once the password and context models were trained, we utilized the password model to identify potential checked-in secrets within our dataset. Subsequently, we employed Soot soot to extract the contextual surroundings of the identified checked-in secrets, specifically capturing six instructions before and six instructions after each occurrence within the same method body, which is the same as the definition of contextual surroundings in the paper of PassFinder passfinder. These surrounding instructions were then used as input to the context model for detection. However, instead of using source code, Soot will output the contents of APK files in Jimple format, an intermediate representation of the Java language. Jimple representations are very similar to Java syntax, so we decided to remain with the Jimple format for the extracted contextual surroundings.
Table 4: The type of checked-in secrets that each tool can detect
@@
The Three-Layer Filter
Amazon, Microsoft, Facebook, Twitter client ID,
Twitter client secret, Twitter access token,
Google API key, Stripe, Square, Picatic,
PayPal Braintree, Twilio, MailGun, MailChimp
LeakScope
Amazon, Microsoft, Facebook, Twitter client ID,
Twitter client secret, Google API key, Stripe,
Square Redhat, IBM cloud
@@
4.1.3 Comparison of the Detection Capability of Each Tool
@@
After making the necessary adaptation for each tool, the detection capability of each tool is listed in Table 4. Overall, we extended LeakScope’s detection capability to align with the Three-Layer Filter. However, it is infeasible to identify a comprehensive set of cloud API signatures for LeakScope that match all the regular expressions used by the Three-Layer Filter and vice versa. This difficulty arises because some cloud providers do not issue secrets that follow specific patterns that can be precisely captured by regular expressions (e.g., Alibaba, Redhat, and IBM). Furthermore, we failed to find the cloud APIs of the cloud providers that the Three-Layer Filter covered (e.g., MailGun, MailChimp, and Picatic). Although the detection capability of LeakScope and the Three-Layer Filter is not entirely equalized, our extension ensured a more meaningful comparison among the tools.
To enable a fair comparison, we present detailed detection results for each cloud provider in the following sections. For the Three-Layer Filter, each detected secret is mapped to its corresponding cloud provider based on the regular expression used for detection. In the case of LeakScope, the mapping is derived from the cloud provider associated with the matched API in our collected cloud API list. For PassFinder, the evaluation dataset includes some secrets that were also detected by the other two tools. In such cases, we adopted the labels generated by those tools as the ground truth. For the remaining cases, we manually mapped the secrets to cloud providers by analyzing the surrounding context. While the detection capabilities of the tools are not perfectly aligned, these mapping strategies allow for a more meaningful and fair comparison. In the following sections, we present the detailed detection results per cloud provider.
4.1.4 Checked-in Secret Validation and Ethical Considerations.
A key challenge in evaluating checked-in secret detection tools is the lack of publicly available ground truth datasets for checked-in secrets in Android apps. To the best of our knowledge, no benchmark dataset currently exists. While full validation of an checked-in secret would ideally involve testing it against the corresponding cloud service, doing so may violate terms of service or raise ethical and legal concerns. To address this, we adopted a manual analysis procedure commonly used in prior checked-in secret research emstudy; secretbench. Specifically, a secret was considered valid if it met the following criteria: (1) it matched the expected format or structure defined by known cloud API key patterns (e.g., prefixes such as AIza for Google), and (2) it appeared in a relevant context, such as within a function call to a known cloud API or near related SDK usage. For instance, in Listing 2, the presence of functions from the Google Cloud Library helped confirm the validity of the suspected checked-in secret.
The manual analysis was conducted independently by two authors. Each author reviewed the format and context of the detected string. In cases of disagreement, both reviewers discussed their reasoning, referenced available documentation when needed, and reached consensus through mutual deliberation. This structured review process allowed us to evaluate suspected checked-in secrets consistently while avoiding direct interaction with external services.
@@
Listing 2: Code snippet with one checked-in secret and its context. It is extracted from an Android app in our experiments.
1GoogleApiClient GoogleClient = x.GoogleClient;
2d.b0.c.l.c(GoogleClient);
3mGoogleApiClient.disconnect();
4GoogleSignInOptions.Builder builder = new Builder;
5builder = builder.requestIdToken("");
6builder = builder.requestEmail();
To complement our manual analysis, we validated the functionality of checked-in secrets identified as Google API keys by sending test requests to the Google Maps service, in order to verify whether the detected secrets were truly usable. We selected this service because the requests do not expose any sensitive user information and allow safe testing without accessing private data. This validation step provides a practical snapshot of the potential real-world impact of leaked secrets by demonstrating that many of them are still active and exploitable. To avoid any significant financial impact, each checked-in secret was used only once. Additionally, for every valid checked-in secret we detected, we notified the developers of the corresponding Android apps.
4.2 RQ2.1 Number of Detected Checked-in Secrets
4.2.1 RQ2.1-Methodology
@@
In this section, we aim to evaluate the effectiveness of the tools by evaluating the number of checked-in secrets that can be detected by each tool. As the detection capability of each tool is not identical, we evaluated the number of detected checked-in secrets identified by each tool across various cloud providers to provide a fairer comparison among the tools. The detailed result of the number of detected checked-in secrets is presented in Table 5. Note that the detected checked-in secrets in this table are not necessarily true positives. In this research question, we demonstrate the overall results reported by each tool. We conduct an analysis on the false positives of the tools in Section 4.3.
@@
Table 5: Number of detected checked-in secrets corresponding to the cloud providers of the three selected tools, the detected checked-in secrets are not verified. NA signifies that the checked-in secrets of the corresponding cloud providers are not in the detection capability of the tool.
@@
The Three-Layer Filter detected the highest number of unique checked-in secrets, while PassFinder failed to detect any. PassFinder’s limited effectiveness can be partially attributed to the context model’s underperformance, with obfuscation being one of the reasons. We further conducted a study to investigate the reasons behind the underperformance of PassFinder. Details of this study are discussed in Section 4.5.
We further categorized all the detected checked-in secrets of each tool to the corresponding cloud providers, as shown in Table 5. Among all the cloud service providers that are supported by each tool, only a few are detected in our dataset. 99.9% of the detected checked-in secrets are from Google Cloud service and Twitter Cloud service. Additionally, LeakScope was able to detect a more significant number of Google checked-in secrets compared to the Three-Layer Filter.
For Twitter Client ID/Secret, the Three-Layer Filter identified 88,172, whereas Leak-Scope only identified 10. This large number of detections from the Three-Layer Filter is expected, as Twitter credentials themselves can take the form as Twitter credentials themselves can take the form as random alphanumeric strings of lengths ranging from 18–25 or 40–50 characters. To capture this property, only relatively loose regular expressions can be used, which unavoidably match a wide range of random strings. Upon analyzing a sample of the 88,172 detected strings, we found that although they exhibit the structural characteristics of API keys, none of them were actual Twitter credentials. We further discuss the implications of this observation in Section 4.3.
@@
4.3 RQ2.2-False Positive
4.3.1 RQ2.2-Methodology
@@
In this section, we aim to compare the false positives reported by each tool. As described in Section 4.1.4, it is impractical and impossible to validate all the obtained checked-in secrets fully. Consequently, we conducted a manual analysis to identify the false positives. Furthermore, since only checked-in secrets belonging to Google API key and Twitter client ID/Secret are detected by at least two tools, we only selected these checked-in secrets for the comparison of false positives. Additionally, PassFinder is excluded from this comparison as it did not detect any checked-in secrets. A further experiment to comprehensively evaluate PassFinder is detailed in Section 4.5.
@@
However, since the Three-Layer Filter detected 88,172 Twitter client IDs/Secrets, manually validating all of these checked-in secrets was unfeasible. To address this, we randomly sampled 383 checked-in secrets from the 88,172, ensuring a 95% confidence level with a 5% margin of error randomsample. Then, we manually analyzed the randomly sampled Twitter client IDs/Secrets and all the Google API keys to identify any false positives as described in Section 4.1.4.
4.3.2 RQ2.2-Result